Show that the following commutation relations follow from the definitions in the previous question: [a+,a−] = −1

[a−,a+] = 1 [H, a+] = ̄hωa+[H, a−] = −̄hωa−
Answer #1

Commutati a Celtnt Laddu 9foeurhore (Part) thame be m ( VRMto Jauw => Ae (a4a)- Caa) Aind (-ipt [,a- and aat a+9 meed we ) CiNorw Gmanimi 2 afa 2 Cm wn) mwh a+ a -1 sho Can we t 2 ma a_ai- r muh Mtes - 2 Mast - 1 aa.-a_a[a, a.1- aat a+a- had Frem Ahe we - 1 aat ata a By ufaslia H= to (a1a- +) hase and we 1 -a at Naw, at Hat a+a at Ha =Au →ン ナー 9+ at9-94 (ar1--4-0-) (-1) a + twat twat twa 4 Ha a-H=a_H aata a HAnd Ha = a aya t9 2 2 6-0-40a)e (1 ummey [at, a-1 - 1 La-, a] 1 (at, HJ [a-, H twar 17 twa (/